Circular No. 9039

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 H1 (LINEAR) E. Guido, G. Sostero, P. Camilleri, and E. Prosperi report their recov-

ery of comet P/2002 LZ_{11} (cf. IAUC 8240) on unfiltered CCD exposures taken remotely with a 0.25-m reflector near Mayhill, NM, U.S.A., on Apr. 17; co-added exposures show an extremely compact coma ∼ 12′′ in diameter and a short tail nearly 25′′ long toward the west. Confirming astrometry on Apr. 18 was obtained by the same team remotely with a 0.37-m reflector near Sonoita, AZ, U.S.A.; co-added exposures show a coma about 8′′ in diameter with a faint extension toward the west.

2009 UT  α_{2000}  δ_{2000}  Mag.
Apr. 17.45321  18^{h}07^{m}32.^{s}66  −17^{o}42′44′′.3  18.8
18.37396  18 07 49.24  −17 38 42.0  19.4

The indicated correction to the predictions on MPC 59599 and in the 2008/2009 Comet Handbook is ΔT = −0.4 day. The following orbital elements by B. G. Marsden are from 154 observations spanning 2002 June 5–2009 Apr. 18 (mean residual 0′′.7). Additional astrometry and orbital elements, residuals, and an ephemeris appear on CBET 1772.

Epoch = 2010 Mar. 25.0 TT

T = 2010 Mar. 5.73233 TT  ω = 107.^{o}76353
e = 0.3532609  Ω = 231.05130 }2000.0
q = 2.3643294 AU  i = 11.52093

a = 3.6557699 AU n^{o} = 0.14100542 P = 6.990 years

COMETS C/2008 Y12–Y15 (SOHO) Further to IAUC 9037, additional near-sun presumed comets have been

found on SOHO website images. C/2008 Y12 and Kreutz sungrazer C/2008 Y15 were stellar in appearance (mag ∼ 7.5). C/2008 Y13, also a Kreutz sungrazer, appeared stellar (mag ∼ 7–7.5) in C3 images, but was very diffuse, faint, and slightly elongated in C2 images. C/2008 Y14 was very diffuse (mag ∼ 8).

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 Y12 Dec. 22.704 17^{h}58.^{m}6 −22^{o}42′ C2  RK 2009-F17
C/2008 Y13  27.888 18 36.1 −25 38 C3/2 AK 2009-F18
C/2008 Y14  27.896 18 20.0 −22 43 C2  JR 2009-F18
C/2008 Y15  29.821 18 45.1 −25 36 C3  BZ 2009-F18
2009 April 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green